OpenStudy (anonymous):

Sandy is working with a carpenter to frame a house. They are using 8-foot-long boards, but each board much be cut to be 94.6 inches long. How much is cut off each board?

OpenStudy (anonymous):

Well first.. how many inches are in a foot? (: 12 right?

OpenStudy (anonymous):

8x12 is 96

OpenStudy (gorv):

1foot =12 inches 8 foot = 8*12=96 inches cuttoff length= 94.6 inches needed to cut =96-94.6=1.4 inches

OpenStudy (gorv):

1.4 inches must be cut from each board